OpenStudy (anonymous):

This is using logarithms. I'm stuck on this problem 6^(x-3/4)=square root of 6 Please Help!

OpenStudy (anonymous):

6^(1/2)=6^(x-3/4) x-3/4=1/2, solve for x

OpenStudy (anonymous):

sqrt(6)=6^(1/2)

OpenStudy (anonymous):

\[\ln 6^{x-3/4}= 6^{1/2}\] \[(x-3/4)\ln 6=(1/2)\ln 6\] \[x-3/4=1/2\] x=5/4

OpenStudy (anonymous):

\[6^{x-3/4}=\sqrt{6}=6^{1/2}\implies x-3/4=1/2 \implies x=5/4\]

OpenStudy (anonymous):

sorry for my ans. I miss ln on the right hand side but the other steps are ok

OpenStudy (anonymous):

OMG Thank Ya'll Sooo Much!!!! I appreciate it! And its fine, its more than what I've done so far.

OpenStudy (anonymous):

:D I am glad that I can give a helping hand to you.

OpenStudy (anonymous):

Well thank you, I appreciate it! :)

OpenStudy (anonymous):

Are you a college student? I am just interested.

OpenStudy (anonymous):

haha yea I am

OpenStudy (anonymous):

me2 and I am from Hong Kong. :D

OpenStudy (anonymous):

Thats awesome, I'm from North America. I bet ya'll are A LOT smarter than we are! lol

OpenStudy (anonymous):

maybe I am just a bit older than you. I am now Form 7, the last year in college. If I am lucky enough, I will be in the university next year.

OpenStudy (anonymous):

Wow thats pretty good. i have a long ways until I get into a university. I'm barely finishing up my first year in a junior college this may.

OpenStudy (anonymous):

oh, i see. Keep going!

OpenStudy (anonymous):

I am! lol Even though its going to take forever to graduate. Good luck to you though!